# Any-stepsize Gradient Descent for Separable Data under Fenchel–Young Losses

### v1

@@ 2 Preliminary on Fenchel–Young losses
Fenchel–Young losses have been introduced by Blondel et al. [9] as a general class of surrogate loss functions for structured prediction and entail the calibration property [52]. We choose Fenchel–Young losses because a vast majority of natural and common loss functions are included in this family. Moreover, the separation margin property, one of the key features of Fenchel–Youngs losses, controls GD behaviors significantly.
@@
Let $\Omega\colon\mathbb{R}^{K}\to\mathbb{R}\cup\left\{{\infty}\right\}$ be a potential function. The induced Fenchel–Young loss $L_{\Omega}\colon\operatorname{\mathrm{dom}}(\Omega^{*})\times\operatorname{\mathrm{dom}}(\Omega)\to\mathbb{R}_{\geq 0}$ by $\Omega$ is defined as
$L_{\Omega}(\bm{\theta};\bm{\mu})\coloneqq\Omega^{*}(\bm{\theta})+\Omega(\bm{\mu})-\left\langle{\bm{\theta}},{\bm{\mu}}\right\rangle.$
In multiclass classification, $L_{\Omega}(\bm{\theta};\bm{\mu})$ measures the proximity of a score vector $\bm{\theta}$ and a target label $\bm{\mu}=\mathbf{e}_{i}$ (for a class $i\in[K]$). By definition, $L_{\Omega}(\cdot,\bm{\mu})$ is convex for any $\bm{\mu}\in\operatorname{\mathrm{dom}}(\Omega)$. Moreover, $L_{\Omega}(\bm{\theta};\bm{\mu})\geq 0$ from the Fenchel–Young inequality, and $L_{\Omega}(\bm{\theta};\bm{\mu})=0$ holds if and only if $\bm{\mu}\in\partial\Omega^{*}(\bm{\theta})$.
We follow Blondel et al. [9, Section 4.4] to instantiate loss functions for binary classification, $K=2$. For simplicity, the following set of assumptions is imposed on a potential function $\Omega$.
@@
$L_{\Omega}(\bm{\theta};\mathbf{e}_{i})=\begin{cases}\phi^{*}(-s)&\text{if $i=1$,}\\ \phi^{*}(s)&\text{if $i=2$,}\end{cases}$
and $\operatorname{\mathrm{dom}}(\phi^{*})=\mathbb{R}$. Hence, the Fenchel–Young loss is simplified as $\phi^{*}(-ys)$ if we relabel two classes $i=1$ and $i=2$ with $y=1$ and $y=-1$, respectively. For the rest of this paper, we suppose a loss function takes the form $\ell(z)\coloneqq\phi^{*}(-z)$ with Assumption 1, which is a symmetric margin-based loss with $z=ys$ [45].
@@
For specific potential functions, Fenchel–Young losses entail a separation margin [9, Section 5], which is a generalized notion of classical margin in support vector machines.
@@
For a loss $\ell\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$, we say $\ell$ has the separation margin property if there exists $m>0$ such that $z\geq m\implies\ell(z)=0$. We call the smallest possible $m$ the separation margin of $\ell$.
Blondel et al. [9, Proposition 6] shows that the existence of the separation margin property can be tested through the subgradient $\partial\phi$.
Proposition 4 (Blondel et al. [9]).
A binary Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption 1 has separation margin if and only if $\partial\phi(\mu)\neq\varnothing$ for any $\mu\in[0,1]$. When $\phi\in\mathcal{C}^{1}((0,1))$ has separation margin $m$, we have
@@
For a differentiable $\phi$, the nonempty-subgradient condition requires that the derivative $\phi^{\prime}(\mu)$ does not explode at the boundary points of the domain $\mu\in\operatorname{\mathrm{dom}}(\phi)=\left\{{0,1}\right\}$. In this case, $\phi$ is not of Legendre-type [42]. As we will see later, the convergence behavior of GD hinges on the separation margin property of a loss function. More detailed analysis of the separation margin property for binary classification can be found in Bao and Sugiyama [6].
@@
If we choose the binary Shannon negentropy $\phi(\mu)=\mu\ln\mu+(1-\mu)\ln(1-\mu)$, we recover the logistic loss $\phi^{*}(-z)=\ln(1+\exp(-z))$. For this reason, (the negative of) a potential function $\phi$ is sometimes referred to as a generalized entropy function. If we choose the negative of the Gini index $\phi(\mu)=\mu^{2}-\mu$, we can generate the modified Huber loss $\phi^{*}(-z)=\max\left\{{0,1-z}\right\}^{2}/4$ if $z\geq-1$ and $\phi^{*}(-z)=-z$ otherwise [58], which is the binarized sparsemax loss [33]. If we choose $\phi(\mu)=\max\left\{{\mu,1-\mu}\right\}$, we recover the hinge loss $\phi^{*}(-z)=\max\left\{{0,1-z}\right\}$. We discuss more examples in Section 4.
@@ 3 Convergence of large stepsize GD under Fenchel–Young losses
We consistently assume the dataset is bounded and linearly separable.
@@
Instead of logistic regression, we choose a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ associated with a binary potential function $\phi$, and minimize the following risk by (GD) with fixed stepsize $\eta>0$ to learn a linear classifier $\mathbf{w}$:
@@
    Fenchel–Young loss. Assume that $\ell(z)$ is a Fenchel–Young loss $\phi^{*}(-z)$ generated by a potential $\phi:\mathbb{R}\to\mathbb{R}\cup\left\{{\infty}\right\}$ such that $\phi\in\mathcal{C}^{2}((0,1))$ satisfies Assumption 1, $\phi$ is strictly convex, and $\phi^{\prime\prime}>0$ on the interval $(0,1)$.
@@
We will later see that $\rho$ characterizes the growth rate of the parameter norm $\|\mathbf{w}_{t}\|$ during GD in (5). Now, we are ready to state our main result, the GD convergence rate for linearly separable data under Fenchel–Young losses. Remarkably, we show convergence without the self-bounding property of a loss function, unlike Wu et al. [55].
@@
Suppose Assumption 2 and consider (GD) with stepsize $\eta>0$ and $\mathbf{w}_{0}=\mathbf{0}$ under a Fenchel–Young loss $\ell$ satisfying Assumption 3. For any $\bar{\varepsilon}\in(0,1)$, let
@@
for $\text{for $\varepsilon\in(0,n\bar{\varepsilon})$,}\quad T>\frac{n^{1+\alpha}}{C_{\phi}\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\varepsilon^{-\alpha}$
holds, then we have $L(\mathbf{w}_{T})\leq\varepsilon$.
This convergence guarantee even applies to non-smooth Fenchel–Young losses as long as Assumption 3 is satisfied—note that $\phi$ must be strongly convex to ensure the smoothness of the associated Fenchel–Young loss [9, Proposition 2.4]. As we see later in Section 4, $\alpha$ and $C_{\phi}$ do neither diverge nor degenerate for arbitrarily small $\bar{\varepsilon}$ under many interesting examples of $\phi$. When $\ell$ has a separation margin, we can simplify the statement of Theorem 5 (by applying Lemma 10 in Appendix A).
@@
Under the same setup with Theorem 5, we additionally assume that $\ell$ has a separation margin $m>0$. If $(\alpha,C_{\phi})$ with (3) satisfies $\alpha,C_{\phi}\in(0,\infty)$ and
for $\text{for $\varepsilon\in(0,n\bar{\varepsilon})$,}\quad T>\frac{n^{1+\alpha}}{C_{\phi}\gamma^{2}}\left(\frac{4m}{\eta}+C_{g}\right)\varepsilon^{-\alpha}$
holds, then we have $L(\mathbf{w}_{T})\leq\varepsilon$.
Therefore, (GD) under many common Fenchel–Young losses converges regardless of the choice of $\eta$. Note that the classical GD convergence analysis under convex smooth functions provides $T=\Omega(\varepsilon^{-1})$. As we see later in Section 4, some loss functions entail better rates with $\alpha<1$, summarized in Table 1.
@@ 3.1 Proof outline
The proof of Theorem 5 essentially relies on the perceptron convergence analysis [38] and the asymptotical order evaluation of rate functions [5]. We sketch the proof in this section to highlight the structure of the GD convergence in our setup and complete the proof in Appendix B.
@@
where $C_{\text{L}}>0$ is a non-degenerate constant independent of $t$. The inequality ()($\clubsuit$)( ♣ ) holds only while perceptron misclassifies some examples. Thus, perceptron correctly classifies all examples after at most $T$ iterations such that $C_{\text{L}}T>C_{\text{U}}(T)$. Such $T$ exists as long as $C_{\text{U}}(t)$ is sublinear in $T$.
When it comes to our setup, an inequality of ()($\clubsuit$)( ♣ )-type is obtained by recursively expanding the update (GD)
$\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle\geq\left\langle{\mathbf{w}_{t-1}},{\mathbf{w}_{*}}\right\rangle+\frac{\gamma\eta}{n}g(\left\langle{\mathbf{w}_{t-1}},{\mathbf{z}_{i_{t-1}}}\right\rangle)\geq\dots\geq\frac{\gamma\eta}{n}\sum_{k=0}^{t-1}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i_{k}}}\right\rangle),$
where $\mathbf{z}_{i_{k}}$ is a misclassified example by $\mathbf{w}_{k}$. Perceptron enjoys an inequality of ()($\clubsuit$)( ♣ )-type immediately because it optimizes the loss function $\ell_{\text{per}}(z)=\max\left\{{-z,0}\right\}$, which yields $g(z)=1$ if $z<0$ (i.e., if misclassified). When considering a Fenchel–Young loss satisfying Assumption 3, we do not have a non-degenerate lower bound for $g(z)$ because we can make $g(z)$ arbitrarily close to zero. Instead, we lower-bound $g(z)$ by a (non-degenerate) error tolerance $\varepsilon_{1}>0$, $g(z)\geq\varepsilon_{1}$, before we attain the $\varepsilon$-optimal loss. Lemma 12 and (a part of) Lemma 14 in Appendix B are relevant to ()($\clubsuit$)( ♣ ).
To obtain an inequality of ()($\diamondsuit$)( ♢ )-type, the standard perceptron analysis directly expands the update (GD) $\|\mathbf{w}_{t}\|^{2}=\|\mathbf{w}_{t-1}-\eta\nabla L(\mathbf{w}_{t-1})\|^{2}$ recursively and upper-bound this by noting that $\ell_{\text{per}}$ has separation margin, leading to $C_{\text{U}}(t)=\mathcal{O}(\sqrt{t})$. Though this is possible for a Fenchel–Young loss with separation margin, we can improve this bound by the split optimization technique, introduced by Wu et al. [55]. Eventually, we can upper-bound $\|\mathbf{w}_{t}\|$ as follows:
@@
In particular, we have $\rho(\lambda)=\mathcal{O}(1)$ when a loss has separation margin (see Lemma 10), and therein $C_{\text{U}}(t)=\mathcal{O}(1)$. This is where separation margin plays a crucial role. We recap the split optimization technique in Lemma 11, based on which Lemma 15 in Appendix C shows this inequality of ()($\diamondsuit$)( ♢ )-type.
The remaining piece is to assess the order of the convergence rate. After solving the inequality (,)($\clubsuit$,$\diamondsuit$)( ♣ , ♢ ) with $t=T$ being the stopping time, we have $T$ as a function of the error tolerance $\varepsilon$, $T=f(\varepsilon)$, where $f$ is a nondecreasing rate function depending on $\phi$. To characterize the asymptotic order at vanishing $\varepsilon$, we attempt to evaluate in the form $f(\varepsilon)\simeq\varepsilon^{\alpha_{0}}$ for an order parameter $\alpha_{0}>0$, which can be estimated by
$\frac{\varepsilon f^{\prime}(\varepsilon)}{f(\varepsilon)}\overset{\varepsilon\downarrow 0}{\longrightarrow}\alpha_{0},\quad\text{if the limit exists.}$.
Thus, the order parameter $\alpha_{0}$ is solely determined by the functional form of potential function $\phi$. This technique has been initially developed in functional analysis to estimate moduli of Banach and Orlicz spaces [46, 22, 11], and recently introduced in convex analysis to approximate a convex function by power functions [23] and estimate moduli of convexity [5, 7]. The general statement of the order evaluation is given in Lemma 13 and instantiated for GD convergence in Lemma 14 in Appendix B.

### v2

@@ 2 Preliminary on Fenchel–Young losses
Fenchel–Young losses have been introduced by Blondel et al. [11] as a general class of surrogate losses for structured prediction, which are classification-calibrated [60]. This can be seen as a Bregman divergence comparing primal and dual points [25, 4]. Despite that the logistic and hinge losses are widely prevailing in practice, we can improve the performance of some prediction tasks by changing specific Fenchel–Young losses, as reported by Roulet et al. [50]. We choose Fenchel–Young losses because a vast majority of convex, Lipschitz, and classification-calibrated losses are included in this class—otherwise, GD convergence is hardly obtained beyond the edge of stability. Moreover, the separation margin property, one of the key features of Fenchel–Young losses, controls GD behaviors significantly.
@@
Let $\Omega\colon\mathbb{R}^{K}\to\mathbb{R}\cup\left\{{\infty}\right\}$ be a potential function. The Fenchel–Young loss $\ell_{\Omega}\colon\operatorname{\mathrm{dom}}(\Omega^{*})\times\operatorname{\mathrm{dom}}(\Omega)\to\mathbb{R}_{\geq 0}$ generated by $\Omega$ is defined as
$\ell_{\Omega}(\bm{\theta};\bm{\mu})\coloneqq\Omega^{*}(\bm{\theta})+\Omega(\bm{\mu})-\left\langle{\bm{\theta}},{\bm{\mu}}\right\rangle.$
In multiclass classification, $\ell_{\Omega}(\bm{\theta};\bm{\mu})$ measures the proximity between a score $\bm{\theta}$ and a target label $\bm{\mu}=\mathbf{e}_{i}$ (for a class $i\in[K]$). By definition, $\ell_{\Omega}(\cdot,\bm{\mu})$ is convex for any $\bm{\mu}\in\operatorname{\mathrm{dom}}(\Omega)$. Moreover, $\ell_{\Omega}(\bm{\theta};\bm{\mu})=0$ holds if and only if $\bm{\mu}\in\partial\Omega^{*}(\bm{\theta})$ due to the equality condition of the Fenchel–Young inequality.
We follow [11, Section 4.4] to consider binary ($K=2$) loss functions. The following set of assumptions is imposed on a potential function $\Omega$. The asymmetric generalization is possible, but we choose to keep the analysis simpler so that we can focus more on the essence of GD convergence.
@@
$\ell_{\Omega}(\bm{\theta};\mathbf{e}_{i})=\begin{cases}\phi^{*}(-s)&\text{if $i=1$,}\\ \phi^{*}(s)&\text{if $i=2$,}\end{cases}$
and $\operatorname{\mathrm{dom}}(\phi^{*})=\mathbb{R}$. Hence, the Fenchel–Young loss is simplified as $\phi^{*}(-ys)$ if we relabel two classes $i=1$ and $i=2$ with $y=1$ and y=1y=-1, respectively. Thus, we suppose the form of a symmetric margin-based loss function $\ell(z)\coloneqq\phi^{*}(-z)$. Therein, a Fenchel–Young loss $\ell$ extends a proper canonical composite loss [47] over the entire prediction space $z\in\mathbb{R}$, as discussed in [7].
@@
For specific potential functions, Fenchel–Young losses entail separation margin [11, Section 5], which is a generalized notion of classical margin in support vector machines.
@@
For a loss $\ell\colon\mathbb{R}\to\mathbb{R}_{\geq 0}$, we say $\ell$ has the separation margin property if there exists $m>0$ such that any prediction $z\geq m$ incurs $\ell(z)=0$. The smallest $m$ is the separation margin of $\ell$.
Hence, $\ell(z)=0$ indicates the prediction $z\in\mathbb{R}$ not only correctly classifies a given point but also has safe margin $m$ away from the classification boundary $z=0$. It is shown that the existence of the separation margin property can be tested through the subgradient $\partial\phi$ [11, Proposition 6].
Proposition 4 ([11]).
A Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ satisfying Assumption˜1 has separation margin if and only if $\partial\phi(\mu)\neq\varnothing$ for any $\mu\in[0,1]$. When $\phi\in\mathcal{C}^{1}((0,1))$ has separation margin $m$,
@@
For a differentiable $\phi$, the nonempty-subgradient condition requires that the derivative $\phi^{\prime}(\mu)$ does not explode at the boundary points of the domain $\mu\in\operatorname{\mathrm{dom}}(\phi)=\left\{{0,1}\right\}$. In this case, $\phi$ is not of Legendre-type [49]. As we will see later, the convergence behavior of GD hinges on the separation margin property of a loss function. More detailed analysis of the separation margin property for binary classification can be found in [7]. In Appendix˜E, we show that a loss satisfying the self-bounding inequality (1) does not have separation margin (but not the other way around).
@@
With the Shannon negentropy $\phi(\mu)=\mu\ln\mu+(1-\mu)\ln(1-\mu)$, we recover the logistic loss $\phi^{*}(-z)=\ln(1+\exp(-z))$. With the negative of the Gini index $\phi(\mu)=\mu^{2}-\mu$, we can generate the modified Huber loss $\phi^{*}(-z)=\max\left\{{0,1-z}\right\}^{2}/4$ if $z\geq-1$ and $\phi^{*}(-z)=-z$ otherwise [67], which is the binarized sparsemax loss [39]. If we choose $\phi(\mu)=\max\left\{{\mu,1-\mu}\right\}$, we recover the hinge loss $\phi^{*}(-z)=\max\left\{{0,1-z}\right\}$. We discuss more examples in Section˜4.
@@ 3 Convergence of large stepsize GD under Fenchel–Young losses
We consistently assume that the dataset is bounded and linearly separable.
@@
Instead of logistic regression, we choose a Fenchel–Young loss $\ell(z)=\phi^{*}(-z)$ associated with a binary potential function $\phi$, and minimize the following risk by (GD) with fixed stepsize $\eta>0$ to learn a linear classifier $\mathbf{w}$: [^2] To generalize the linear model, one straightforward way is to focus on deep homogeneous networks [56]. We stay on the linear model for now because the straightforward extension to deep homogeneous networks may not significantly change the problem structure.
@@
    Fenchel–Young loss. Assume that $\ell(z)$ is a Fenchel–Young loss $\phi^{*}(-z)$ generated by a potential $\phi:\mathbb{R}\to\mathbb{R}\cup\left\{{\infty}\right\}$ such that $\phi\in\mathcal{C}^{2}((0,1))$ satisfies Assumption˜1, $\phi$ is strictly convex, and $\phi^{\prime\prime}>0$ on the interval $(0,1)$.
@@
We will later see that $\rho$ characterizes the growth rate of the parameter norm $\|\mathbf{w}_{t}\|$ during GD in (5). This notion is inherited from [63]. Now, we are ready to state our main result, the GD rate to attain arbitrarily small loss for linearly separable data under Fenchel–Young losses. Remarkably, we show convergence without the self-bounding property of a loss function, unlike [63].
@@
Suppose Assumption˜2 and consider (GD) with stepsize $\eta>0$ and $\mathbf{w}_{0}=\mathbf{0}$ under a Fenchel–Young loss $\ell$ satisfying Assumption˜3. For any $\bar{\varepsilon}\in(0,1)$, let
@@
$\text{for $\varepsilon\in(0,\bar{\varepsilon})$,}\quad T>\frac{n}{C_{\phi}\gamma^{2}}\left(\frac{4\sqrt{\rho(\gamma^{2}\eta T)}}{\eta}+C_{g}\right)\varepsilon^{-\alpha}$
holds, then we have $\min_{t\in[T]}L(\mathbf{w}_{t})\leq\varepsilon$.
This convergence guarantee even applies to non-smooth Fenchel–Young losses as long as Assumption˜3 is satisfied—note that $\phi$ must be strongly convex to ensure the smoothness of the associated Fenchel–Young loss [11, Proposition 2.4]. As seen later in Section˜4, $\alpha$ and $C_{\phi}$ neither diverge nor degenerate for arbitrarily small $\bar{\varepsilon}$ under many examples of $\phi$. When $\ell$ has separation margin, Lemma˜9 in Appendix˜A provides a finite upper bound on $\rho$, yielding the following simpler form.
@@
Under the same setup with Theorem˜5, we additionally assume that $\ell$ has separation margin $m>0$. For any $\bar{\varepsilon}\in(0,1)$, if $(\alpha,C_{\phi})$ defined in (3) satisfies $\alpha,C_{\phi}\in(0,\infty)$ and
$\text{for $\varepsilon\in(0,\bar{\varepsilon})$,}\quad T>\frac{n}{C_{\phi}\gamma^{2}}\left(\frac{4m}{\eta}+C_{g}\right)\varepsilon^{-\alpha}$
holds, then we have $\min_{t\in[T]}L(\mathbf{w}_{t})\leq\varepsilon$.
A loss function without separation margin does not have finite $\rho$, which typically yields slower convergence as we see in Section˜4. Therefore, (GD) operated on many common Fenchel–Young losses converges under the separability, regardless of the choice of $\eta$. Note that the classical GD convergence analysis under convex smooth functions provides $T=\Omega(\varepsilon^{-1})$. As we see later in Section˜4, some loss functions entail better rates with $\alpha<1$, summarized in Table˜1.
@@ 3.1 Proof outline
The proof of Theorem˜5 essentially relies on the perceptron convergence analysis [44] and the asymptotical order evaluation of rate functions [6]. We sketch the proof in this section to highlight the structure of the GD convergence in our setup and complete the proof in Appendix˜B.
@@
where $C_{\text{L}}>0$ is a non-degenerate constant independent of $t$. The inequality $(\clubsuit)$ holds only while perceptron misclassifies some examples. Thus, perceptron correctly classifies all examples after at most $T$ iterations such that $C_{\text{L}}T>C_{\text{U}}(T)$. Such $T$ exists as long as $C_{\text{U}}(t)$ is sublinear in $T$.
When it comes to our setup, an inequality $(\clubsuit)$ is obtained by recursively expanding the update (GD) [^3] This expansion relies on $\left\langle{\mathbf{w}_{*}},{\mathbf{z}_{i_{t-1}}}\right\rangle\geq\gamma$, namely, the linear separability in Assumption 2. To lift the linear separability assumption, we may require to introduce additional distributional assumptions here.
$\left\langle{\mathbf{w}_{t}},{\mathbf{w}_{*}}\right\rangle\geq\left\langle{\mathbf{w}_{t-1}},{\mathbf{w}_{*}}\right\rangle+\frac{\gamma\eta}{n}g(\left\langle{\mathbf{w}_{t-1}},{\mathbf{z}_{i_{t-1}}}\right\rangle)\geq\dots\geq\frac{\gamma\eta}{n}\sum_{k=0}^{t-1}g(\left\langle{\mathbf{w}_{k}},{\mathbf{z}_{i_{k}}}\right\rangle)\eqqcolon\gamma\eta\widetilde{G}(\mathbf{w}_{k}),$
where $\mathbf{z}_{i_{k}}$ is a misclassified example by $\mathbf{w}_{k}$. Perceptron enjoys an inequality of $(\clubsuit)$-type immediately because it optimizes the loss function $\ell_{\text{per}}(z)=\max\left\{{-z,0}\right\}$, which yields $g(z)=1$ if $z<0$ (i.e., if misclassified). When considering a Fenchel–Young loss satisfying Assumption˜3, we do not have a non-degenerate lower bound for $g(z)$ because we can make $g(z)$ arbitrarily close to zero. Instead, we lower-bound $g(z)$ by a (non-degenerate) error tolerance $\varepsilon_{1}>0$, $g(z)\geq\varepsilon_{1}$, before we attain the $\varepsilon$-optimal loss. Lemma˜11 and (a part of) Lemma˜13 in Appendix˜B are relevant to $(\clubsuit)$. Note that the perceptron argument is used in [63] but in a different way: they control $L(\mathbf{w}_{k})$ through the upper bound on $\widetilde{G}(\mathbf{w}_{k})$ (see Lemma˜25). This is applicable only to self-bounding losses.
To obtain an inequality of $(\diamondsuit)$-type, by following the standard perceptron analysis, we directly expand the update (GD) $\|\mathbf{w}_{t}\|^{2}=\|\mathbf{w}_{t-1}-\eta\nabla L(\mathbf{w}_{t-1})\|^{2}$ recursively, and upper-bound it by noting that $\ell_{\text{per}}$ has separation margin, leading to $C_{\text{U}}(t)=\mathcal{O}(\sqrt{t})$. Though this is possible for a Fenchel–Young loss with separation margin, we can improve this bound by borrowing the split optimization technique, introduced by [63]. Eventually, we can upper-bound $\|\mathbf{w}_{t}\|$ as follows:
@@
In particular, we have $\rho(\lambda)=\mathcal{O}(1)$ when a loss has separation margin (see Lemma˜9), and therein $C_{\text{U}}(t)=\mathcal{O}(1)$. This is where separation margin plays a crucial role. We recap the split optimization technique in Lemma˜10, based on which Lemma˜19 in Appendix˜D shows this inequality of $(\diamondsuit)$-type.
The remaining piece is to assess the order of the convergence rate. After solving the inequality $(\clubsuit,\diamondsuit)$ with $t=T$ being the stopping time, we have $T$ as a function of the error tolerance $\varepsilon$, $T=f(\varepsilon)$, where $f$ is a nondecreasing rate function depending on $\phi$. To characterize the asymptotic order at vanishing $\varepsilon$, we attempt to evaluate in the form $f(\varepsilon)\simeq\varepsilon^{\alpha_{0}}$ for an order parameter $\alpha_{0}>0$, which can be estimated by
$\frac{\varepsilon f^{\prime}(\varepsilon)}{f(\varepsilon)}\overset{\varepsilon\downarrow 0}{\longrightarrow}\alpha_{0},\quad\text{if the limit exists.}$
Thus, the order parameter $\alpha_{0}$ is solely determined by the functional form of potential function $\phi$. This technique has been initially developed in functional analysis to estimate moduli of Banach and Orlicz spaces [53, 27, 13], and recently introduced in convex analysis to approximate a convex function by power functions [28] and estimate moduli of convexity [6, 8]. The general statement of the order evaluation is given in Lemma˜12 and instantiated for GD convergence in Lemma˜13 in Appendix˜B.